# It’s Not Easy Being Green: On the Energy Efficiency of Programming Languages

### v3

@@ Abstract
Does the choice of programming language affect energy consumption? Previous highly visible studies have established associations between certain programming languages and energy consumption. A causal misinterpretation of this work has led academics and industry leaders to use or support certain languages based on their claimed impact on energy consumption. This paper tackles this causal question directly. It first corrects and improves the measurement methodology used by prior work. It then develops a detailed causal model capturing the complex relationship between programming language choice and energy consumption. This model identifies and incorporates several critical but previously overlooked factors that affect energy usage. These factors, such as distinguishing programming languages from their implementations, the impact of the application implementations themselves, the number of active cores, and memory activity, can significantly skew energy consumption measurements if not accounted for. We show—via empirical experiments, improved methodology, and careful examination of anomalies—that when these factors are controlled for, notable discrepancies in prior work vanish. Our analysis suggests that the choice of programming language implementation has no significant impact on energy consumption beyond execution time.
@@ I Introduction
These studies have received wide attention, both in academic and industrial circles. They have been collectively cited over 730 times per Google Scholar (as of May 2025). The results—especially the ranking of programming languages—have had an unusually visible impact in industry, and are routinely quoted on social networks and in blog posts. As Figure 2 illustrates, the rankings have been cited by executives and engineers from Amazon [12, 13, 14, 15], Intel [16], SAP [17], and other companies [18] to argue for business decisions and to advocate for a shift in programming languages with an eye towards sustainability. Often, these rankings have been harnessed to support the adoption of Rust, which ranks as one of the most energy-efficient languages while providing safety guarantees that languages like C and C++ lack.
@@
(b) Keynote talk at KubeCon + CloudNativeCon Europe 2022: “Coding with Python over Rust for large scale applications can mean a difference of up to 75×75$\times$75 × in energy usage”.
@@
The above analysis and approach suffer from numerous other methodological flaws (§II-B, §III, §IV-B, §IV-E). However, the primary focus of this paper is carefully addressing the question: does the choice of programming language affect energy consumption? We embrace this question by developing a rich causal model of the relationship between programming language and energy consumption. Figure 1 presents a causal diagram representing our final model. Causal diagrams illustrate how different factors influence each other, with arrows showing the direction of these influences [19]. We build this model incrementally, incorporating the various factors at play and their relationships, and provide quantitative and/or qualitative evidence supporting each step.
@@
The remainder of this paper is organized as follows: Section II introduces and presents a critical analysis of prior work. Section III describes the improved methodology we adopt in our experiments, detailing the measurement tool and the benchmarks used. Section IV incrementally builds the causal model of the relationship between programming language and energy consumption, explaining and correcting anomalies found in prior work at each step. Section VI concludes with a discussion of other related work.
@@ II-A Overview
TABLE I: Partial results from Pereira et al. (§II-A), showing average execution time and energy consumption for the languages discussed in this paper, normalized to C. Boldface denotes anomalous results that Sections II-B, IV-B, and IV-E address.
@@ II-B Critique
Programming Language versus Implementation: Programming languages define the syntax and semantics, but it is their implementations that primarily influence performance. While some languages have only a single, widely-used implementation such as Rust or Go, others have multiple implementations, each with their own performance characteristics. For instance, Pereira et al. treat Ruby and JRuby as different languages, while they are in fact two separate implementations of the same Ruby language. The papers uses different benchmark implementations to compare these two Ruby implementations, confounding their comparison.
@@
Apparent Anomalies in Results: Some results reported by Pereira et al. are counter-intuitive, and are presented without investigation or explanation. C++ is reported as being 34% less energy efficient and 56% slower than C. Since C++ is approximately a superset of C, and both share the same compiler, optimizations, and code generation backend, we would expect identical energy consumption and execution time. Similarly, TypeScript is reported as being 4.8×4.8$\times$4.8 × less energy efficient and 7.1×7.1$\times$7.1 × slower than JavaScript. TypeScript is a strict superset of JavaScript: any valid JavaScript program is a valid TypeScript program. The same Node.js runtime system is used for both languages. The compilation process for TypeScript may insert or modify code to support older JavaScript standards, but we do not expect this to result in any significant performance overhead. Java, C#, and Go are reported as 1.89×1.89$\times$1.89 ×, 3.14×3.14$\times$3.14 ×, and 2.83×2.83$\times$2.83 × slower than C, respectively. These numbers are unexpectedly high as these implementations are known to be highly optimized [21]. We expect low overhead from garbage collection costs, plus initial just-in-time compilation overhead for Java and C#. Lua and TypeScript both stand out as exhibiting significantly lower normalized energy consumption than normalized execution time; this is explained by the fact that nearly all of the Lua and TypeScript benchmark implementations are sequential, while the implementations of these benchmarks in other languages are parallelized. Section IV-D1 characterizes the impact of the number of active cores on energy consumption.
@@
Total memory usage is measured with Python’s memory_profiler, which samples the RSS at a frequency of ten times per second. This metric is directly proportional to execution time, which is itself directly proportional to energy consumption.
@@
Failure to Control for Language Implementation Specifics: Most language implementations have an initial cost to import and set up necessary in-memory data structures. Language implementations using a just-in-time compiler also require an initial warmup period [23, 24]. These start-up costs are amplified by the benchmarks’ short execution times, which in some cases run for less than a fraction of a second. For example, measuring only the first iteration of a short-lived Java benchmark may not be indicative of Java’s overall performance. Section IV-B1 shows that this first iteration can be up to 3×3$\times$3 × slower than subsequent ones. Garbage collection also can have a significant impact on both execution time and memory usage [25], and can be fine-tuned to obtain better performance.
III Methodology
TABLE II: Versions of compilers and runtime systems used (§III-A). All are most recent as of September 2024.
Implementation Language(s) Version
LLVM / Clang C, C++ 19
Rust Rust 1.81.0
OpenJDK Java 21.0.4+7
Go Go 1.23.1
C# / .NET C# 8.0.8
Node.js JavaScript, TypeScript 20.17.0
PHP PHP 8.3.11
tsc TypeScript 5.6.2
CPython Python 3.12.6
PyPy Python 7.3.17
Lua Lua 5.4.7
LuaJIT Lua 87ae18a
III-A Setup
We conduct experiments in this paper using a server equipped with two Intel Xeon Gold 6430 processors totaling 128 logical cores and 128GB of memory, running Linux version 6.8.0-45-generic. Table II details specific compiler and runtime system versions. Experiments in both this paper and Pereira et al.’s use a server architecture.
Benchmark runs are isolated in Docker containers to ensure reproducibility. Docker introduces negligible performance overhead [26], which we confirm for the experiments detailed in this paper by obtaining equivalent results with and without Docker. Docker also provides easy access to control groups, making it possible to place restrictions on CPU usage or to pin programs to execute on specific cores. Section IV-C details how we eliminate the effect of varying concurrency in different benchmark implementations by limiting benchmarks to execute on a single core.
III-B Measurement Tool
@@
III-B1 RAPL
Starting with the Sandy Bridge microarchitecture (2011), most Intel processors provide a power management interface called Running Average Power Limit (RAPL). RAPL allows measurement of the energy consumption of various parts of the system [28]. The interface is exposed via Model Specific Registers (MSRs), of which the following measure energy consumption:
@@ III-B1 RAPL
III-B2 Performance Counters
Hardware performance counters are a feature of modern processors that track various events, such as total cycles, cache misses, branch misses, and so on. They are often used to analyze program performance, notably when profiling. These counters are 64 bits in length and thus not susceptible to overflow, so it suffices to read them once at the beginning and once at the end of each experiment.
Memory Activity
Section IV-D2 highlights the impact of memory activity on energy consumption: a majority of memory energy consumption is dependent on the rate of read and write operations. We monitor the last level cache (LLC) using performance counters to estimate memory activity. Each LLC hit means that the cache already had a copy of the data, and therefore no further memory activity occurs. On the other hand, each LLC miss means that the system must fetch data from memory. Therefore, the number of LLC misses is a good proxy for memory activity.
Average CPU Usage
There are multiple ways to measure average core usage while a program is running. For simplicity, we leverage performance counters provided by the operating system. The Linux kernel exposes the task-clock software event counter, which aggregates in nanoseconds the time spent on all processor cores. To obtain average core usage, we divide this counter’s value by the total execution time of the program.
III-C Benchmarks
This paper primarily focuses its evaluation on the benchmarks used in the original Pereira et al. study, namely the Computer Language Benchmark Game. In some analyses, we extend that evaluation with two additional benchmark suites: SPEC CPU 2017 [30] and DaCapo Chopin [31]. The following subsections provide a brief overview of each of these benchmark suites.
III-C1 Computer Language Benchmark Game
TABLE III: Descriptions of benchmarks from the Computer Language Benchmark Game (§III-C1), along with the range of lines of code (LoC) across languages for each benchmark as reported by cloc.
@@ III-C1 Computer Language Benchmark Game
The Computer Language Benchmark Game [20] (CLBG) is a collection of small programs implemented in many different programming languages (22 to 287 lines of code). We use the same benchmark implementations as Pereira et al. [27], which is in effect a snapshot of the fastest versions of the benchmarks as available on the CLBG repository at the time of the original work. Table III provides a brief description of each benchmark. We limit our analysis to eleven languages across thirteen implementations: C, C#, C++, Go, Java, JavaScript, Lua (Lua and LuaJIT), PHP, Python (CPython and PyPy), Rust, and TypeScript. These languages comprise 11 of the top 12 most popular general programming languages in the 2024 Stack Overflow survey [32] (excluding Kotlin).
@@
III-C2 SPEC CPU 2017
TABLE IV: SPEC CPU 2017 benchmarks (§III-C2).
@@ III-C2 SPEC CPU 2017
TABLE V: DaCapo benchmarks (§III-C3).
@@
SPEC CPU 2017 [30] is a widely used benchmark suite for evaluating systems on C, C++ and Fortran code. The suite consists of large real-world programs and workloads, ranging between 1k and 1.3M lines of code. We use the latest available version (1.1.9) and run all benchmarks in the “speed” category using the reference input set. We omit a single benchmark, cam4, as it did not successfully run on our platform. Table IV briefly describes of each benchmark.
III-C3 DaCapo
DaCapo [31] is a widely used benchmark suite for Java applications. The suite consists of large real-world programs and workloads, ranging between 25k and 6M lines of code. We use the latest available version (Chopin) and run all benchmarks using the default input set. Table V provides a brief description of each benchmark.
IV Causal Model for Energy Consumption in Programming Languages
This section builds piece by piece the causal model presented as a diagram in its final form in Figure 1. Causal diagrams illustrate how different factors influence each other, with arrows indicating cause-to-effect relationships [19]. We additionally use gray arrows to represent comparatively weaker relationships, as detailed in the following sections. Each subsection below gradually adds nodes and edges to the diagram in order to obtain a rich causal model that captures the essential factors in the relationship between the choice of programming language and energy consumption. Figures 3, 4, 7, and 8 build upon each other, with new elements in orange. They culminate in the final model represented in Figure 1.
IV-A Starting Point
ProgrammingLanguageEnergy
IV-B Programming Languages and Implementations
ProgrammingLanguageProgrammingLanguageImplementationEnergy
Figure 4 introduces the important distinction between the programming languages themselves and their implementations. Some languages have a single well-known implementation, or a blurry line between language and implementation, such as Rust or Go. Others have multiple implementations, such as C/C++ with GCC, Clang, or MSVC, Java with various JVMs, Python with CPython and PyPy, or Lua with LuaJIT. Finally, there can be significant differences between specific versions of the same implementation. Our model treats those different versions as different implementations altogether. For example, recent CPython releases each successively delivered significant performance improvements, notably with the 3.11 release achieving a 25% runtime speedup over 3.10 [33].
Of course, programming languages have an effect on implementation possibilities. They may dictate the general memory layout of objects, the need for a garbage collector, or the possibility of ahead-of-time compilation. For example, Java forces object dereferences to be indirect, and has no alternative for memory reclamation other than a garbage collector. Other dynamic features such as dynamic typing, reflection, or eval further limit the range of options available to implementers.

### v4

@@ Abstract
Does the choice of programming language affect energy consumption? Previous highly visible studies have established associations between certain programming languages and energy consumption. A causal misinterpretation of this work has led academics and industry leaders to use or support certain languages based on their claimed impact on energy consumption. This paper tackles this causal question directly: it develops a detailed causal model capturing the complex relationship between programming language choice and energy consumption. This model identifies and incorporates several critical but previously overlooked factors that affect energy usage. These factors, such as distinguishing programming languages from their implementations, the impact of the application implementations themselves, the number of active cores, and memory activity, can significantly skew energy consumption measurements if not accounted for. We show—via empirical experiments, improved methodology, and careful examination of anomalies—that when these factors are controlled for, notable discrepancies in prior work vanish. Our analysis suggests that the choice of programming language implementation has no significant impact on energy consumption beyond execution time.
@@ I Introduction
These studies have received wide attention, both in academic and industrial circles. They have been collectively cited over 800 times per Google Scholar (as of October 2025). The results—especially the ranking of programming languages—have had an unusually visible impact in industry, and are routinely quoted on social networks and in blog posts. As Figure 2 illustrates, the rankings have been cited by executives and engineers from Amazon [12, 13, 14, 15], Intel [16], SAP [17], and other companies [18] to argue for business decisions and to advocate for a shift in programming languages with an eye towards sustainability. Often, these rankings have been harnessed to support the adoption of Rust, which ranks as one of the most energy-efficient languages while providing safety guarantees that languages like C and C++ lack.
@@
(b) Keynote talk at KubeCon + CloudNativeCon Europe 2022: “Coding with Python over Rust for large scale applications can mean a difference of up to $75\times$ in energy usage”.
@@
The above analysis and approach suffer from numerous other methodological flaws (§II-B, §V, §VI-C). However, the primary focus of this paper is carefully addressing the question: does the choice of programming language affect energy consumption? We embrace this question by developing a rich causal model of the relationship between programming language and energy consumption. Figure 1 presents a causal diagram representing our final model. Causal diagrams illustrate how different factors influence each other, with arrows showing the direction of these influences [19]. We build this model incrementally, incorporating the various factors at play and their relationships, and provide quantitative and/or qualitative evidence supporting each step.
@@ II-A Overview
TABLE I: Partial results from Pereira et al. (§II-A), showing average execution time and energy consumption for the languages discussed in this paper, normalized to C. Boldface denotes anomalous results that Sections II-B, VI-C address.
@@ II-B Critique
Programming Language versus Implementation: Programming languages define the syntax and semantics, but it is their implementations that primarily influence performance. While some languages have only a single, widely-used implementation such as Rust or Go, others have multiple implementations, each with their own performance characteristics. For instance, Pereira et al. treat Ruby and JRuby as different languages, while they are in fact two separate implementations of the same Ruby language. The papers use different benchmark implementations to compare these two Ruby implementations, confounding their comparison.
@@
Apparent Anomalies in Results: Some results reported by Pereira et al. are counter-intuitive, and are presented without investigation or explanation. C++ is reported as being 34% less energy efficient and 56% slower than C. Since C++ is approximately a superset of C, and both share the same compiler, optimizations, and code generation backend, we would expect identical energy consumption and execution time. Similarly, TypeScript is reported as being $4.8\times$ less energy efficient and $7.1\times$ slower than JavaScript. TypeScript is a strict superset of JavaScript: any valid JavaScript program is a valid TypeScript program. The same Node.js runtime system is used for both languages. The compilation process for TypeScript may insert or modify code to support older JavaScript standards, but we do not expect this to result in any significant performance overhead. Java, C#, and Go are reported as $1.89\times$, $3.14\times$, and $2.83\times$ slower than C, respectively. These numbers are unexpectedly high as these implementations are known to be highly optimized [21]. We expect low overhead from garbage collection costs, plus initial just-in-time compilation overhead for Java and C#. Lua and TypeScript both stand out as exhibiting significantly lower normalized energy consumption than normalized execution time; this is explained by the fact that nearly all of the Lua and TypeScript benchmark implementations are sequential, while the implementations of these benchmarks in other languages are parallelized. Section VI-B characterizes the impact of the number of active cores on energy consumption.
@@
The total memory usage is measured by using Python’s memory_profiler, which samples the RSS at a frequency of ten times per second. This metric is directly proportional to execution time, which is itself directly proportional to energy consumption.
@@
Failure to Control for Language Implementation Characteristics: Most language implementations have an initial cost to import and set up necessary in-memory data structures. Language implementations using a just-in-time compiler also require an initial warmup period [23, 24]. These start-up costs are amplified by the benchmarks’ short execution times, which in some cases run for less than a fraction of a second. For example, measuring only the first iteration of a short-lived Java benchmark may not be indicative of Java’s overall performance. Section VI-C shows that this first iteration can be up to $3\times$ slower than subsequent ones. Garbage collection also can have a significant impact on both execution time and memory usage [25], and can be fine-tuned to obtain better performance.
III Research Questions
This paper provides numerous improvements to methodology and results over prior work. Section IV first introduces and gradually builds a causal model capturing all factors standing between choice of programming language and energy consumption. Section V describes the improved methodology used in this paper, including correction of technical errors and enhanced data gathering with performance counters. Finally, using both the causal model and improved methodology, Section VI addresses the following research questions:
  • RQ1: Do some language implementations consume more energy than others?
  • RQ2: What are the key contributors to power draw standing between choice of programming language and energy consumption?
  • RQ3: Can observed anomalies in prior work be explained through the lens of our causal model?
IV Causal Model for Energy Consumption in Programming Languages
This section builds piece by piece the causal model presented as a diagram in its final form in Figure 1. Causal diagrams illustrate how different factors influence each other, with directed edges indicating cause-to-effect relationships [19]. Formally, edges indicate that changes in the origin node impact the probability distribution of its descendants. Our diagrams additionally use gray arrows to represent comparatively weaker relationships, as detailed in the following sections. Each subsection below gradually adds nodes and edges to the diagram in order to obtain a rich causal model that captures the essential factors in the relationship between the choice of programming language and energy consumption. Figures 3, 4, 5, and 6 build upon each other, with new elements in orange. They culminate in the final model represented in Figure 1.
IV-A Starting Point
ProgrammingLanguageEnergy
IV-B Programming Languages and Implementations
ProgrammingLanguageProgrammingLanguageImplementationEnergy
Figure 4 introduces the important distinction between the programming languages themselves and their implementations. Some languages have a single well-known implementation, or a blurry line between language and implementation, such as Rust or Go. Others have multiple implementations, such as C/C++ with GCC, Clang, or MSVC, Java with various JVMs, Python with CPython and PyPy, or Lua with LuaJIT. Finally, there can be significant differences between specific versions of the same implementation. Our model treats those different versions as different implementations altogether. For example, recent CPython releases each successively delivered significant performance improvements, notably with the 3.11 release achieving a 25% speedup over 3.10 [26].
Of course, programming languages have an effect on implementation possibilities. They may dictate the general memory layout of objects, the need for a garbage collector, or the possibility of ahead-of-time compilation. For example, Java has no alternative for memory reclamation other than a garbage collector. Other dynamic features such as dynamic typing, reflection, or eval further limit the range of options available to implementers.
The impact of the characteristics of programming languages on their implementations is transitively included in our model: the language has a causal influence on the implementation, which in turn has a direct causal impact on execution time. The performance and energy consumption impact of those language features is mediated by the implementation. Dynamic dispatch, for example, may be implemented or optimized with v-tables, fat pointers, monomorphic and polymorphic inline caches, or profile-guided optimization.
IV-C Decomposing Energy into Power and Time
Fundamentally, energy consumption is the product of power and time. Figure 5 decomposes these two factors. Section VI-B demonstrates that power draw is constant across all languages, implementations, and benchmarks. Claims of the contrary in prior work were in large part due to missing multicore accounting, directly comparing sequential and concurrent benchmarks.
In theory, runtime systems running in parallel to the user program may affect energy consumption beyond variance in execution time, since increased core usage will increase power draw, as Section VI-B details below. This hidden increased parallelism could occur with a parallel garbage collector or just-in-time compilation threads executing alongside the main program. In practice, compilation threads are primarily active only during startup, and both are rarely significant compared to the main program’s execution and other general overhead costs. For this reason, we draw this edge in gray in our model.
IV-D Contributors to Power Draw
Previous research has demonstrated that processor and memory energy usage are the main variable contributors to power draw [27]. We add these two factors to our model in Figure 6. For the studied programs on our experimental platform, the ratio of memory energy consumption to CPU energy consumption remains between 2 and 8% (Section VI-B). Further, memory activity is not something that is easily controlled by the programmer. While certain algorithms or data structures maximize cache locality, there is eventually a limit stemming from the LLC’s fixed, relatively small size. Programming languages or their implementations may impose additional overhead with a garbage collector, memory allocation overhead, or object memory layout. Because of the difficulty of controlling memory activity, as well as its comparatively smaller impact compared to CPU activity on power draw, we draw this edge in gray in our model.
IV-E Application Implementations
We add a final node to our diagram: the application implementations themselves, which have the most direct impact on energy consumption. Figure 1 presents the final causal model developed in this paper. Application implementation decisions dictate the level of parallelism, the algorithms used, and the data structures employed.
IV-F Other Factors
V Enhanced Methodology
Our methodology improves upon prior work in several ways. First, it corrects technical errors in the measurement tool used by Pereira et al. [29]. It also enhances data gathering by collecting performance counter readings, which we use to identify key contributors to energy consumption. Finally, it expands the set of benchmarks used to include two additional widely-used benchmark suites.
V-A Measurement Tool
@@
RAPL: Starting with the Sandy Bridge microarchitecture (2011), most Intel processors provide a power management interface called Running Average Power Limit (RAPL). RAPL allows measurement of the energy consumption of various parts of the system [30]. The interface is exposed via Model Specific Registers (MSRs), of which the following measure energy consumption:
@@
Performance Counters: Hardware performance counters are a feature of modern processors that track various events, such as total cycles, cache misses, branch misses, and so on. They are often used to analyze program performance, notably when profiling. These counters are 64 bits in length and thus not susceptible to overflow, so it suffices to read them once at the beginning and once at the end of each experiment.
Memory Activity: Section IV-D highlights the impact of memory activity on energy consumption: a majority of memory energy consumption is dependent on the rate of read and write operations. We monitor the last level cache (LLC) using performance counters to estimate memory activity. Each LLC hit means that the cache already had a copy of the data, and therefore no further memory activity occurs. On the other hand, each LLC miss means that the system must fetch data from memory. Therefore, the number of LLC misses is a good proxy for memory activity.
Average CPU Usage: There are multiple ways to measure average core usage while a program is running. For simplicity, we leverage performance counters provided by the operating system. The Linux kernel exposes the task-clock software event counter, which aggregates in nanoseconds the time spent on all processor cores. To obtain average core usage, we divide this counter’s value by the total execution time of the program.
V-B Benchmarks
This paper extends the evaluation performed by the original Pereira et al. study by incorporating two additional large-scale benchmark suites. We first include all possible benchmark implementations from the Computer Language Benchmark Game used by Pereira et al. that successfully compile and run (118 benchmarks). Using the same set of benchmarks allows reproduction and re-examination of results in Section VI-C. In addition, the analysis in Sections VI-A and VI-B includes two more widely-used benchmark suites: SPEC CPU 2017 [32] (18 benchmarks) and DaCapo Chopin [33] (22 benchmarks). The following subsections provide a brief overview of each of these benchmark suites.
TABLE II: Computer Language Benchmark Game benchmarks (§V-B), along with the range of lines of code (LoC) across languages for each benchmark as reported by cloc.
@@ V-B Benchmarks
Computer Language Benchmark Game: The Computer Language Benchmark Game [20] (CLBG) is a collection of small programs implemented in many different programming languages (22 to 287 lines of code). We use the same benchmark implementations as Pereira et al. [29], which is in effect a snapshot of the fastest versions of the benchmarks as available on the CLBG repository at the time of the original work. Table II provides a brief description of each benchmark. We limit our analysis to eleven languages across thirteen implementations: C, C#, C++, Go, Java, JavaScript, Lua (Lua and LuaJIT), PHP, Python (CPython and PyPy), Rust, and TypeScript. These languages comprise 11 of the top 12 most popular general programming languages in the 2024 Stack Overflow survey [34] (excluding Kotlin).
@@
TABLE III: SPEC CPU 2017 benchmarks (§V-B).
@@
TABLE IV: DaCapo benchmarks (§V-B).
@@
SPEC CPU 2017: SPEC CPU 2017 [32] is a widely used benchmark suite for evaluating systems on C, C++ and Fortran code. The suite consists of large real-world programs and workloads, ranging between 1k and 1.3M lines of code. We use the latest available version (1.1.9) and run all benchmarks in the “speed” category using the reference input set. We omit a single benchmark, cam4, as it did not successfully run on our platform. Table III briefly describes each benchmark.
DaCapo: DaCapo [33] is a widely used benchmark suite for Java applications. The suite consists of large real-world programs and workloads, ranging between 25k and 6M lines of code. We use the latest available version (Chopin) and run all benchmarks using the default input set. Table IV provides a brief description of each benchmark.
VI Evaluation
This section addresses the research questions posed in Section III, using the causal model and enhanced methodology introduced in Sections IV and V.
TABLE V: Versions of compilers and runtime systems used (§VI). All are most recent as of September 2024.
Implementation Language(s) Version
LLVM / Clang C, C++ 19
Rust Rust 1.81.0
OpenJDK Java 21.0.4+7
Go Go 1.23.1
C# / .NET C# 8.0.8
Node.js JavaScript, TypeScript 20.17.0
PHP PHP 8.3.11
tsc TypeScript 5.6.2
CPython Python 3.12.6
PyPy Python 7.3.17
Lua Lua 5.4.7
LuaJIT Lua 87ae18a
We conduct experiments in this paper using a server equipped with two Intel Xeon Gold 6430 processors totaling 128 logical cores and 128GB of memory, running Linux version 6.8.0-45-generic. Table V details specific compiler and runtime system versions. Benchmark runs are isolated in Docker containers to ensure reproducibility. Docker introduces negligible performance overhead [35], which we confirm for the experiments detailed in this paper by obtaining equivalent results without Docker. Docker provides easy access to control groups, making it possible to place restrictions on CPU usage or to pin programs to execute on specific cores (Section VI-B).
VI-A RQ1: Do Some Language Implementations Consume More Energy Than Others?